# Oracle-Based Multistep Strategy for Solving Polynomial Systems Over Finite Fields and Algebraic Cryptanalysis of the Aradi Cipher

### v3

@@ Abstract.
In this paper, we present a new recursive formulation of the corresponding algorithm based on a Depth-First Search strategy, along with a novel complexity analysis leveraging tree structures. We also introduce the notion of an “oracle function”, which is intended to determine whether evaluating a new variable is required to simplify the current polynomial system. This notion allows us to unify all previously proposed variants of the multistep strategy, including the classical hybrid approach, by appropriately selecting the oracle function.
@@ 3. Complexity
Let us analyze the complexity of MultiSolve when computing $V_{\mathbb{F}}(J(i_{1},\ldots,i_{k}))$. Recall that for $k=0$, we have $J()=J$. Indeed, we assume that the computation time of GBElimLin is negligible, due to the low value of the parameter $d_{1}$. Likewise, we assume that invoking the function Oracle is also negligible.
@@
By assuming that the oracle function is accurate, the complexity of MultiSolve is reduced to the number of computed tame cases. Indeed, the oracle function associated with the MultiSolve algorithm in [14] is assumed to be accurate. Moreover, given an integer $B>0$, the standard hybrid strategy, as described in [2, 3], can be obtained as the MultiSolve algorithm implementing an accurate oracle function of type
@@
Note that in the above result, we consider the total number of wild cases, including those that are computed by GroebnerSafe, which incur some cost. If the function Oracle is accurate, such as OracleNRV in the attack on Trivium [14], there is no such cost. However, if this oracle is not applicable, for instance in the sparse case where the number of remaining variables is insufficient to accurately predict the output status of GroebnerSafe computations, different features of the input polynomial systems has to be considered to design an accurate oracle. Developing such oracle functions is generally a challenging task, which we plan to pursue in future work with the aid of Machine Learning (ML) techniques.
@@
By applying Theorem 3.3, one obtains the following result.
@@
$\frac{N}{L}=\frac{qL-1}{L(q-1)}=1+\frac{1}{q-1}(1-\frac{1}{L})$
In particular, for $q=2$, we have N/L=21/LN/L=2-1/L, which implies $1\leq N/L<2$.
@@
where $L$ is the number of computed tame cases. This $L$ also represents the total number of GroebnerSafe executions when using a perfect oracle. Computing the ratio $N/L$ yields the formula of the claim. ∎
Note that we may also define a constant wild function, which reduces MultiSolve to the recursive application of GBElimLin. For $d_{1}=0$, this corresponds to brute-force solving by evaluating all possible values of the variables in the base finite field $\mathbb{F}$}.
@@
it is necessary to apply MultiSolve to all ideals $J(i_{1},\ldots,i_{k})$. The classical hybrid strategy consists of using the oracle function ${\textsc{OracleH}}_{C}$. If a perfect or, at least, an accurate oracle function is not available, a first approximation could be given by ${\textsc{OracleH}}_{B}$. Note that using this oracle actually corresponds to executing MultiSolve with the function OracleT, for all ideals $J(i_{1},\ldots,i_{k},i_{k+1},\ldots,i_{k+B})$ where $1\leq i_{k+1},\ldots,i_{k+B}\leq q$. A thorough analysis of the distributions of computed wild and tame cases, possibly supported by ML techniques, could further help refine the oracle function.
@@ 7. Conclusions and further directions
As a further direction for this line of research, we suggest developing oracle functions based on Machine Learning or Deep Learning algorithms, trained on the ideals, and corresponding Macaulay matrices, generated by MultiSolve. A preliminary experimental study is currently underway. Another promising direction lies in the development of a parallel implementation of the MultiSolve algorithm, as well as in its extension to polynomial systems admitting multiple solutions, which could be leveraged in discrete optimization tasks.

### v4

@@ Abstract.
In this paper, we present a new formulation of the corresponding algorithm based on a Depth-First Search strategy, along with a novel complexity analysis leveraging tree structures. We also introduce the notion of an “oracle function”, which is intended to determine whether evaluating a new variable is required to simplify the current polynomial system. This notion allows us to unify all previously proposed variants of the multistep strategy, including the classical hybrid approach, by appropriately selecting the oracle function.
@@ 3. Complexity
Let us analyze the complexity of MultiSolve when computing $V_{\mathbb{F}}(J(i_{1},\ldots,i_{k}))$. Recall that for $k=0$, we have $J()=J$. Indeed, we assume that the computation time of GBElimLin is negligible, due to the low value of the parameter $d_{1}$. Likewise, we assume that invoking the Oracle function is negligible, as these functions merely involve comparing numerical data or performing a forward pass of a neural network, operations that are typically very fast.
@@
By assuming that the oracle function is accurate, the complexity of MultiSolve is reduced to the number of computed tame cases. In other words, there are no computed wild cases incurring the timeout cost $\tau$, and consequently no significative overhead arises in dealing with wild and tame cases according to the multistep solving strategy. Indeed, the oracle function associated with the MultiSolve algorithm in [14] is assumed to be accurate. Moreover, given an integer $B>0$, the standard hybrid strategy, as described in [2, 3], can be obtained as the MultiSolve algorithm implementing an accurate oracle function of type
@@
When the oracle function is accurate, such as OracleNRV in the attack on Trivium [14], all computational effort is devoted, up to negligible overhead, to solving the tame cases, and the above result implies the superiority of the multistep strategy over the hybrid one. However, if OracleNRV is not applicable, for instance in the sparse case where the number of remaining variables is insufficient to accurately predict the output status of GroebnerSafe computations, different features of the input polynomial systems has to be considered to design an accurate oracle.
A promising approach we are currently pursuing is the use of machine learning techniques to build a data-driven oracle for predicting the wall-clock runtime of the Block-Wiedemann procedure when applied to sparse Macaulay matrices arising in the multistep process. We adopt a permutation-invariant Deep Sets architecture [19], which processes each row independently, aggregates the resulting embeddings to enforce invariance with respect to row ordering, and combines them with coarse-grained metadata. The resulting global representation is then passed to a final regression head that outputs a scalar runtime prediction. The model is trained with mean squared error loss and evaluated through repeated cross-validation, assessing both predictive accuracy and how runtime relates to sparsity, matrix dimensions, and construction parameters. These results will be reported in a forthcoming paper.
@@
We now compare the worst-case and optimal oracles, showing that the overhead introduced by computed wild cases remains moderate, especially when compared with the high complexity levels encountered in cryptography. Applying Theorem 3.3 then yields the following result.
@@
$\frac{N}{L}=1+\frac{1-1/L}{q-1}$
and therefore
$1\leq\frac{N}{L}<\frac{q}{q-1}.$
In particular, for $q=2$, we have N/L=21/LN/L=2-1/L and $1\leq N/L<2$.
@@
where $L$ is the number of computed tame cases. This $L$ also represents the total number of GroebnerSafe executions when using a perfect oracle. Computing the ratio $N/L$ yields
$\frac{N}{L}=\frac{qL-1}{L(q-1)}=1+\frac{1-1/L}{q-1}$
To establish the bounds, observe that the function
$f(L)=1+\frac{1-1/L}{q-1}$
is strictly increasing for $L\geq 1$, since $f^{\prime}(L)=\frac{1}{(q-1)L^{2}}>0$. Hence, the minimum of $f$ over $L\geq 1$ is attained at $L=1$, giving
$\min_{L\geq 1}\frac{N}{L}=f(1)=1.$
Moreover, since 11/L<11-1/L<1 for all finite $L>1$, we obtain the strict upper bound
$\frac{N}{L}=f(L)<1+\frac{1}{q-1}=\frac{q}{q-1}.$
Since each GroebnerSafe call, whether corresponding to a tame or a wild case, is bounded by time $\tau$, the total running time using OracleT is at most a factor of $\frac{q}{q-1}$ larger than with a perfect oracle. In particular, for $q=2$ this overhead is strictly less than $2$, corresponding to at most a one-bit increase in the exponent of the complexity, while for large field sizes $q$ it becomes negligible as $N/L\to 1$.
Note that choosing a constant wild oracle function reduces MultiSolve to a recursive application of GBElimLin. For $d_{1}=0$, this corresponds to brute-force enumeration of all variable assignments over the base finite field $\mathbb{F}$}.
@@
it is necessary to apply MultiSolve to all ideals $J(i_{1},\ldots,i_{k})$. The classical hybrid strategy consists of using the oracle function ${\textsc{OracleH}}_{C}$. If a perfect or, at least, an accurate oracle function is not available, a first approximation can be provided by ${\textsc{OracleH}}_{B}$. Note that using this oracle actually corresponds to executing MultiSolve with the function OracleT, for all ideals $J(i_{1},\ldots,i_{k},i_{k+1},\ldots,i_{k+B})$ where $1\leq i_{k+1},\ldots,i_{k+B}\leq q$. A more refined oracle function may be derived from a detailed analysis of the distribution of wild and tame cases, for example using machine learning techniques.
@@ 7. Conclusions and further directions
As a further direction for this line of research, we suggest developing oracle functions based on machine learning algorithms, trained on the ideals and corresponding Macaulay matrices generated by MultiSolve. A preliminary experimental study is currently underway. Another promising direction lies in the development of a parallel implementation of the MultiSolve algorithm, as well as in its extension to polynomial systems admitting multiple solutions, which could be leveraged in discrete optimization tasks.
Acknowledgments
The authors would like to thank the anonymous reviewers for their careful reading of the manuscript and for their constructive comments, which enhanced the clarity and readability of the work.